OpenStudy (youngnephew34):

What is the perimeter of triangle ABC with the vertices A(-2,9), B(7,-3),C(-2,-3) in the coordinate plan A)21 units B) 15 units C)34 units D)36 units PLZZZZZZZZZZ HELPPPPPPP MEDALS AND FAN

OpenStudy (mrnood):

First sketch the triangle on a coordinate plane For example  and put the values of x & y for each point

OpenStudy (anonymous):

36

OpenStudy (mrnood):

I think she is wrong

OpenStudy (anonymous):

9-(-3)= 12 7-(-2)= 9 12x12= 144 9x9= 81 144+81= 225 square root of 225= 15 so 12+9+15= 36

OpenStudy (mrnood):

*actually I have no idea if she is correct - I just want the poster to work it out - he cannot trust someone giving a direct answer without explanation - so I sow the seed of doubt....

OpenStudy (youngnephew34):

thx guys